\begin{thebibliography}{25}
\providecommand{\natexlab}[1]{#1}
\providecommand{\url}[1]{\texttt{#1}}
\expandafter\ifx\csname urlstyle\endcsname\relax
  \providecommand{\doi}[1]{doi: #1}\else
  \providecommand{\doi}{doi: \begingroup \urlstyle{rm}\Url}\fi

\bibitem[Bai et~al.(2018)Bai, Kolter, and Koltun]{bai_kolter_koltun_2018}
Shaojie Bai, J.~Zico Kolter, and Vladlen Koltun.
\newblock An empirical evaluation of generic convolutional and recurrent
  networks for sequence modeling, 2018.
\newblock URL \url{https://arxiv.org/abs/1803.01271}.

\bibitem[Boyd and Chua(1985)]{boyd85fading}
Stephen Boyd and Leon~O. Chua.
\newblock Fading memory and the problem of approximating nonlinear operators
  with {Volterra} series.
\newblock \emph{IEEE Transactions on Circuits and Systems}, {CAS-32}\penalty0
  (11):\penalty0 1150--1161, 1985.

\bibitem[Chelba et~al.(2017)Chelba, Norouzi, and
  Bengio]{chelba_norouzi_bengio_2017}
Ciprian Chelba, Mohammad Norouzi, and Samy Bengio.
\newblock N-gram language modeling using recurrent neural network estimation,
  2017.
\newblock URL \url{https://arxiv.org/abs/1703.10724}.

\bibitem[Dauphin et~al.(2017)Dauphin, Fan, Auli, and
  Grangier]{dauphin_fan_auli_grangier_2016}
Yann~N. Dauphin, Angela Fan, Michael Auli, and David Grangier.
\newblock Language modeling with gated convolutional networks.
\newblock In \emph{International Conference on Machine Learning}, 2017.

\bibitem[{DeVore} and Lorentz(1993)]{devore1993apx}
Ronald~A. {DeVore} and George~G. Lorentz.
\newblock \emph{Constructive Approximation}.
\newblock Springer-Verlag, Berlin, 1993.

\bibitem[Gehring et~al.(2017)Gehring, Auli, Grangier, Yarats, and
  Dauphin]{gehring_et_al_2017}
Jonas Gehring, Michael Auli, David Grangier, Denis Yarats, and Yann~N. Dauphin.
\newblock Convolutional sequence to sequence learning.
\newblock In \emph{International Conference on Machine Learning}, 2017.

\bibitem[Hanin and Sellke(2018)]{hanin18minwidth}
Boris Hanin and Mark Sellke.
\newblock Approximating continuous functions by {ReLU} nets of minimal width,
  2018.
\newblock URL \url{http://arxiv.org/abs/1710.11278}.

\bibitem[Johnson and Zhang(2017)]{johnson_zhang_2017}
Rie Johnson and Tong Zhang.
\newblock Deep pyramid convolutional neural networks for text categorization.
\newblock In \emph{Proceedings of the 55th Annual Meeting of the Association
  for Computational Linguistics (Volume 1: Long Papers)}, pages 562--570,
  Vancouver, Canada, July 2017. Association for Computational Linguistics.
\newblock \doi{10.18653/v1/P17-1052}.
\newblock URL \url{https://www.aclweb.org/anthology/P17-1052}.

\bibitem[Kalchbrenner et~al.(2016)Kalchbrenner, Espeholt, Simonyan, van~den
  Oord, Graves, and Kavukcuoglu]{kalchbrenner_et_al_2016}
Nal Kalchbrenner, Lasse Espeholt, Karen Simonyan, Aaron van~den Oord, Alex
  Graves, and Koray Kavukcuoglu.
\newblock Neural machine translation in linear time, 2016.
\newblock URL \url{https://arxiv.org/abs/1610.10099}.

\bibitem[Kim and Braatz(2014)]{kim14lure}
{Kwang-Ki}~K. Kim and Richard~D. Braatz.
\newblock Observer-based output feedback control of discrete-time {Lur'e}
  systems with sector-bounded slope-restricted nonlinearities.
\newblock \emph{International Journal of Robust and Nonlinear Control},
  24:\penalty0 2458--2472, 2014.

\bibitem[Liang and Srikant(2017)]{liang2017deep}
Shiyu Liang and R.~Srikant.
\newblock Why deep neural networks for function approximation?
\newblock In \emph{International Conference on Learning Representations}, 2017.

\bibitem[Miller and Hardt(2019)]{miller2019stable}
John Miller and Moritz Hardt.
\newblock Stable recurrent models.
\newblock In \emph{International Conference on Learning Representations}, 2019.

\bibitem[Park and Sandberg(1992)]{park1992criteria}
Jooyoung Park and Irwin~W. Sandberg.
\newblock Criteria for the approximation of nonlinear systems.
\newblock \emph{IEEE Transactions on Circuits and Systems --- {I: Fundamental
  Theory and Applications}}, 39\penalty0 (8):\penalty0 673--676, 1992.

\bibitem[Pavlov et~al.(2006)Pavlov, {van de Wouw}, and
  Nijmeijer]{pavlov06nonlinea}
Alexey Pavlov, Nathan {van de Wouw}, and Henk Nijmeijer.
\newblock \emph{Uniform Output Regulation of Nonlinear Systems: A Convergent
  Dynamics Approach}.
\newblock Birkh\"auser, 2006.

\bibitem[Sandberg(1991)]{sandberg91structure}
Irwin~W. Sandberg.
\newblock Structure theorems for nonlinear systems.
\newblock \emph{Multidimensional Systems and Signal Processing}, 2:\penalty0
  267--286, 1991.

\bibitem[Sandberg and Xu(1993)]{sandberg93circle}
Irwin~W. Sandberg and Lilian~Y. Xu.
\newblock Steady-state errors in discrete-time control systems.
\newblock \emph{Automatica}, 29\penalty0 (2):\penalty0 523--526, 1993.

\bibitem[Sarkans and Logemann(2016)]{sarkans2016lure}
Elvijs Sarkans and Hartmut Logemann.
\newblock Input-to-state stability of discrete-time {Lur'e} systems.
\newblock \emph{SIAM Journal on Control and Optimization}, 54\penalty0
  (3):\penalty0 1739--1768, 2016.

\bibitem[Sharan et~al.(2018)Sharan, Kakade, Liang, and
  Valiant]{sharan_et_al_2016}
Vatsal Sharan, Sham Kakade, Percy Liang, and Gregory Valiant.
\newblock Prediction with a short memory.
\newblock In \emph{Symposium on Theory of Computing}, 2018.

\bibitem[Sontag(1998)]{sontag98control}
Eduardo~D. Sontag.
\newblock \emph{Mathematical Control Theory: Deterministic finite Dimensional
  Systems}.
\newblock Springer-Verlag, 1998.

\bibitem[Tran et~al.(2017)Tran, R\"uffler, and Kellett]{tran2017convergent}
Duc~N. Tran, Bj\"orn~S. R\"uffler, and Christopher~M. Kellett.
\newblock Convergence properties for discrete-time nonlinear systems.
\newblock \emph{IEEE Transactions on Automatic Control}, 2017.

\bibitem[Tsypkin(1964)]{tsypkin64circle}
Yakov~Z. Tsypkin.
\newblock A criterion of absolute stability for sampled-data systems with
  monotone characteristics of the nonlinear element.
\newblock \emph{Doklady Akademii Nauk {SSSR}}, 155\penalty0 (5):\penalty0
  1029--1032, 1964.
\newblock In Russian.

\bibitem[Vaidyanathan(1985)]{vaidyanathan85dtbr}
Palghat~P. Vaidyanathan.
\newblock The discrete-time bounded-real lemma in digital filtering.
\newblock \emph{IEEE Transactions on Circuits and Systems}, {CAS-32}\penalty0
  (9):\penalty0 918--924, September 1985.

\bibitem[{van den Oord} et~al.(2016){van den Oord}, Dieleman, Zen, Simonyan,
  Vinyals, Graves, Kalchbrenner, Senior, and Kavukcuoglu]{oord_et_al_2016}
Aaron {van den Oord}, Sander Dieleman, Heiga Zen, Karen Simonyan, Oriol
  Vinyals, Alex Graves, Nal Kalchbrenner, Andrew Senior, and Koray Kavukcuoglu.
\newblock Wavenet: A generative model for raw audio, 2016.
\newblock URL \url{https://arxiv.org/abs/1609.03499}.

\bibitem[Wu et~al.(2016)Wu, Schuster, Chen, Le, Norouzi, Macherey, Krikun, Cao,
  Gao, Macherey, Klingner, Shah, Johnson, Liu, Kaiser, Gouws, Kato, Kudo,
  Kazawa, Stevens, Kurian, Patil, Wang, Young, Smith, Riesa, Rudnick, Vinyals,
  Corrado, Hughes, and Dean]{wu_et_al_2016}
Yonghui Wu, Mike Schuster, Zhifeng Chen, Quoc~V. Le, Mohammad Norouzi, Wolfgang
  Macherey, Maxim Krikun, Yuan Cao, Qin Gao, Klaus Macherey, Jeff Klingner,
  Apurva Shah, Melvin Johnson, Xiaobing Liu, {\L}ukasz Kaiser, Stephan Gouws,
  Yoshikiyo Kato, Taku Kudo, Hideto Kazawa, Keith Stevens, George Kurian,
  Nishant Patil, Wei Wang, Cliff Young, Jason Smith, Jason Riesa, Alex Rudnick,
  Oriol Vinyals, Greg Corrado, Macduff Hughes, and Jeffrey Dean.
\newblock Google's neural machine translation system: Bridging the gap between
  human and machine translation, 2016.

\bibitem[Yin et~al.(2017)Yin, Kann, Yu, and
  Sch{\"u}tze]{yin_kann_yu_schutze_2017}
Wenpeng Yin, Katharina Kann, Mo~Yu, and Hinrich Sch{\"u}tze.
\newblock Comparative study of {CNN} and {RNN} for natural language processing,
  2017.
\newblock URL \url{https://arxiv.org/abs/1702.01923}.

\end{thebibliography}

\documentclass{article}

\usepackage[final]{neurips_2019}

\usepackage[utf8]{inputenc} 
\usepackage[T1]{fontenc}    
\usepackage{hyperref}       
\usepackage{url}            
\usepackage{booktabs}       
\usepackage{amsfonts}       
\usepackage{nicefrac}       
\usepackage{microtype}      

\usepackage{amsmath,amssymb,amsthm}
\usepackage{commath}
\usepackage{verbatim}

\usepackage[mathcal]{euscript}

    \def\ddefloop#1{\ifx\ddefloop#1\else\ddef{#1}\expandafter\ddefloop\fi}

    \def\ddef#1{\expandafter\def\csname c#1\endcsname{\ensuremath{\mathcal{#1}}}}
    \ddefloop ABCDEFGHIJKLMNOPQRSTUVWXYZ\ddefloop

    \def\ddef#1{\expandafter\def\csname s#1\endcsname{\ensuremath{\mathsf{#1}}}}
    \ddefloop ABCDEFGHIJKLMNOPQRSTUVWXYZ\ddefloop
	
    \def\ddef#1{\expandafter\def\csname bb#1\endcsname{\ensuremath{\mathbb{#1}}}}
    \ddefloop ABCDEFGHIJKLMNOPQRSTUVWXYZ\ddefloop
	
    \def\ddef#1{\expandafter\def\csname bd#1\endcsname{\ensuremath{\boldsymbol{#1}}}}
    \ddefloop ABCDEFGHIJKLMNOPQRSTUVWXYZ\ddefloop
	
    \def\ddef#1{\expandafter\def\csname bd#1\endcsname{\ensuremath{\boldsymbol{#1}}}}
    \ddefloop abcdefghijklmnopqrstuvwxyz\ddefloop

    \def\Reals{\mathbb{R}}

    \def\deq{:=}

    \def\wh#1{\widehat{#1}}
    \def\bd#1{\boldsymbol{#1}}

    \def\1{{\mathbf 1}}

    \def\eps{\varepsilon}
	\def\ReLU{\operatorname{ReLU}}

    \newtheorem{theorem}{Theorem}[section]
    \newtheorem{definition}[theorem]{Definition}
    
    \newtheorem{proposition}[theorem]{Proposition}
    \newtheorem{corollary}[theorem]{Corollary}
    \newtheorem{assumption}[theorem]{Assumption}
    \newtheorem{remark}[theorem]{Remark}

\title{Universal Approximation of Input-Output Maps by Temporal Convolutional Nets}

\author{%
  Joshua Hanson
  \\
  University of Illinois\\
  Urbana, IL 61801 \\
  \texttt{jmh4@illinois.edu} \\
  \And
  Maxim Raginsky \\
  University of Illinois \\
 Urbana, IL 61801 \\
  \texttt{maxim@illinois.edu} \\
}

\begin{document}

\maketitle

\begin{abstract}
There has been a recent shift in sequence-to-sequence modeling from recurrent network architectures  to convolutional network architectures due to computational advantages in training and operation while still achieving competitive performance. For systems having limited long-term temporal dependencies, the approximation capability of recurrent networks is essentially equivalent to that of temporal convolutional nets (TCNs). We prove that TCNs can approximate a large class of input-output maps having approximately finite memory to arbitrary error tolerance. Furthermore, we derive quantitative approximation rates for deep ReLU TCNs in terms of the width and depth of the network and modulus of continuity of the original input-output map, and apply these results to input-output maps of systems that admit finite-dimensional state-space realizations (i.e., recurrent models).
\end{abstract}

	\section{Introduction}
	
Until recently, recurrent networks have been considered the de facto standard for modeling input-output maps that transform sequences to sequences. Convolutional network architectures are becoming favorable alternatives for several applications due to reduced computational overhead incurred during both training and regular operation, while often performing as well as or better than recurrent architectures in practice. The computational advantage of convolutional networks follows from the lack of feedback elements, which enables shifted copies of the input sequence to be processed in parallel rather than sequentially \citep{gehring_et_al_2017}. Convolutional architectures have demonstrated exceptional accuracy in sequence modeling tasks that have typically been approached using recurrent architectures, such as machine translation, audio generation, and language modeling \citep{dauphin_fan_auli_grangier_2016,kalchbrenner_et_al_2016,oord_et_al_2016,wu_et_al_2016,gehring_et_al_2017,johnson_zhang_2017}.

	One explanation for this shift is that both convolutional and recurrent architectures are inherently suited to modeling systems with limited long-term dependencies. Recurrent models possess infinite memory (the output at each time is a function of the initial conditions and the entire history of inputs until that time), and thus are strictly more expressive than finite-memory autoregressive models. However, in synthetic stress tests designed to measure the ability to model long-term behavior, recurrent architectures often fail to learn long sequences \citep{bai_kolter_koltun_2018}. Furthermore, this unlimited memory property is usually unnecessary, which is supported in theory \citep{sharan_et_al_2016} and in practice \citep{chelba_norouzi_bengio_2017,gehring_et_al_2017}. In situations where it is only important to learn finite-length sequences, feedforward architectures based on temporal convolutions (temporal convolutional nets, or TCNs) can achieve similar results and even outperform recurrent nets \citep{dauphin_fan_auli_grangier_2016,yin_kann_yu_schutze_2017,bai_kolter_koltun_2018}.

These results prompt a closer look at the conditions under which convolutional architectures provide better approximation than recurrent architectures. Recent work by \cite{miller2019stable} has shown that recurrent models that are exponentially stable (in the sense that the effect of the initial conditions on the output decays exponentially with time) can be efficiently approximated by feedforward models.  A key consequence is that exponentially stable recurrent models can be approximated by systems that only consider a finite number of recent values of the input sequence for determining the value of the subsequent output.

However, this notion of stability is inherently tied to a particular state-space realization, and it is not difficult to come up with examples of sequence-to-sequence maps that have both a stable and an unstable state-space realization (e.g., simply by adding unstable states that do not affect the output). This suggests that the question of approximating sequence-to-sequence maps by feedforward convolutional maps should be studied by abstracting away the notion of stability and only requiring that the system output depend appreciably on recent input values and negligibly on input values in the distant past. The formalization of this property was introduced by \cite{sandberg91structure} under the name of \textit{approximately finite memory}, building on earlier work by \cite{boyd85fading}. Outputs of systems characterized by this property can be approximated by the output of the same system when applied to a truncated version of the input sequence. These systems are naturally suited to be modeled using TCNs, which by construction only operate on values of the input sequence for times within a finite horizon into the past.

	In this work, we aim to develop quantitative results for the approximation capability of TCNs for modeling input-output maps that have the properties of causality, time invariance, and approximately finite memory. In Section~\ref{sec:afm}, we introduce the necessary definitions and review the approximately finite memory property due to \cite{sandberg91structure}. Section~\ref{sec:apx} gives the main result for approximating input-output maps by ReLU TCNs, together with a quantitative result on the equivalence between approximately finite memory and a related notion of \textit{fading memory} \citep{boyd85fading,park1992criteria}. These results are applied in Section~\ref{sec:recurrent} to  recurrent models that are \textit{incrementally stable} \citep{tran2017convergent}, i.e., the influence of the initial condition is asymptotically negligible. We show that incrementally stable recurrent models have approximately finite memory, and then use this formalism to derive a generalization of the result of \cite{miller2019stable}. We provide a comparison in Section~\ref{sec:comparison} to other architectures used for approximating input-output maps. All omitted proofs are provided in the Supplementary Material.
	
	\section{Input-output maps and approximately finite memory}
	\label{sec:afm}

Let $\cS$ denote the set of all real-valued sequences $\bdu = (u_t)_{t\in\bbZ_+}$, where $\bbZ_+ \deq \{0,1,2,\ldots\}$. An \textit{input-output map} (or i/o map, for short) is a nonlinear operator $\sF : \cS \to \cS$ that maps an input sequence $\bdu \in \cS$ to an output sequence $\bdy = \sF\bdu \in \cS$. (We are considering real-valued input and output sequences for simplicity; all our results carry over to vector-valued sequences at the expense of additional notation.) We will denote the application and the composition of i/o maps by concatenation. In this paper, we are concerned with i/o maps $\sF$ that are:
\begin{itemize}
	\item \textit{causal} --- for any $t \in \bbZ_+$, $\bdu_{0:t} = \bdv_{0:t}$ implies $(\sF\bdu)_t = (\sF\bdv)_t$, where $\bdu_{0:t} \deq (u_0,\ldots,u_t)$;
	\item \textit{time-invariant} --- for any $k \in \bbZ_+$,
    \begin{equation*}
        (\sF \sR^k \bdu)_t =
        \begin{cases}
            (\sF \bdu)_{t-k}, & \text{for } t \geq k \\
            0, & \text{for } 0 \leq t < k
        \end{cases},
    \end{equation*}
    where $\sR : \cS \to \cS$ is the right shift operator $(\sR \bdu)_t \deq u_{t-1}\1_{\{t \ge 1\}}$.
\end{itemize}
The key notion we will work with is that of \textit{approximately finite memory} \citep{sandberg91structure}:
\begin{definition} An i/o map $\sF$ has {\em approximately finite memory} on a set of inputs $\cM \subseteq \cS$ if for any $\eps > 0$ there exists $m \in \bbZ_+$, such that
	\begin{align}\label{eq:AFM}
		\sup_{\bdu \in \cM} \sup_{t \in \bbZ_+}  \left|(\sF \bdu)_t - (\sF \sW_{t,m}\bdu)_t\right| \leq \eps,
	\end{align}
	where $\sW_{t,m} : \cS \to \cS$ is the {\em windowing operator} $(\sW_{t,m}\bdu)_\tau \deq u_\tau \1_{\{\max\{t-m,0\} \le \tau \le t\}}$. We will denote by $m^*_\sF(\eps)$ the smallest $m \in \bbZ_+$, for which \eqref{eq:AFM} holds.
\end{definition}
If $m^*_\sF(0) < \infty$, then we say that $\sF$ has \textit{finite memory} on $\cM$. If $\sF$ is causal and time-invariant, this is equivalent to the existence of an integer $m \in \bbZ_+$ and a nonlinear functional $f : \Reals^{m+1} \to \Reals$, such that $f(0,\ldots,0) = 0$ and, for any $\bdu \in \cM$ and any $t \in \bbZ_+$,
\begin{align}\label{eq:FM}
	(\sF\bdu)_t = f(u_{t-m},u_{t-m+1},\ldots,u_t),
\end{align}	
with the convention that $u_s = 0$ if $s < 0$. In this work, we will focus on the important case when $f$ is a feedforward neural net with rectified linear unit (ReLU) activations $\ReLU(x) \deq \max\{x,0\}$. That is, there exist $k$ affine maps $A_i : \Reals^{d_i} \to \Reals^{d_{i+1}}$ with $d_1 = m+1$ and $d_{k+1} = 1$, such that $f$ is given by the composition
\begin{align*}
f = A_k \circ \ReLU \circ A_{k-1} \circ \ReLU \circ \ldots \circ \ReLU \circ A_1,
\end{align*}
where, for any $r \ge 1$, $\ReLU(x_1,\ldots,x_r) \deq (\ReLU(x_1),\ldots,\ReLU(x_r))$. Here, $k$ is the depth (number of layers) and $\max\{d_2,\ldots,d_k\}$ is the width (largest number of units in any hidden layer).

\begin{definition} An i/o map $\sF$ is a {\em ReLU temporal convolutional net} (or ReLU TCN, for short) with context length $m$ if \eqref{eq:FM} holds for some feedforward ReLU neural net $f : \Reals^{m+1} \to \Reals$. 
\end{definition}

\begin{remark} {\em While such an $\sF$ is evidently causal, it is generally not time-invariant unless $f(0,\ldots,0)=0$.}
\end{remark}

\section{The universal approximation theorem}
\label{sec:apx}

In this section, we state and prove our main result: any causal and time-invariant i/o map that has approximately finite memory and satisfies an additional continuity condition can be approximated arbitrarily well by a ReLU temporal convolutional net. In what follows, we will consider i/o maps with uniformly bounded inputs, i.e., inputs in the set 
\begin{align*}
\cM(R) \deq	\{ \bdu \in \cS : \|\bdu\|_\infty \deq \sup_{t \in \bbZ_+}|u_t| \le R\} \qquad \text{for some $R > 0$}.
\end{align*}
For any $t \in \bbZ_+$ and any $\bdu \in \cM(R)$, the finite subsequence $\bdu_{0:t} = (u_0,\ldots,u_t)$ is an element of the cube $[-R,R]^{t+1} \subset \Reals^{t+1}$; conversely, any vector $\bdx \in [-R,R]^{t+1}$ can be embedded into $\cM(R)$ by setting $u_s = x_s \1_{\{0 \le s \le t\}}$. To any causal and time-invariant i/o map $\sF$ we can associate the nonlinear functional $\tilde{\sF}_t : \Reals^{t+1} \to \Reals$ defined in the obvious way: for any $\bdx = (x_0,x_1,\ldots,x_t) \in \Reals^{t+1}$,
\begin{align*}
	\tilde{\sF}_t(\bdx) \deq (\sF \bdu)_t,
\end{align*}
where $\bdu \in \cS$ is any input such that $u_s = x_s$ for $s \in \{0,1,\ldots,t\}$ (the values of $u_s$ for $s > t$ can be arbitrary by causality). We impose the following assumptions on $\sF$:

\begin{assumption}\label{as:AFM} The i/o map $\sF$ has approximately finite memory on $\cM(R)$.
\end{assumption}
\begin{assumption}\label{as:cont} For any $t \in \bbZ_+$, the functional $\tilde{\sF}_t : \Reals^{t+1} \to \Reals$ is uniformly continuous on $[-R,R]^{t+1}$ with  modulus of continuity
		\begin{align*}
			\omega_{t,\sF}(\delta) \deq \sup \left\{ |\tilde{\sF}_t(\bdx)-\tilde{\sF}_t(\bdx')| : \bdx,\bdx' \in [-R,R]^{t+1}, \|\bdx - \bdx'\|_\infty \le \delta \right\},
		\end{align*}
		and inverse modulus of continuity
		\begin{align*}
			\omega^{-1}_{t,\sF}(\eps) \deq \sup \left\{ \delta > 0 : \omega_{t,\sF}(\delta) \le \eps\right\}.
		\end{align*}
		where $\|\bdx\|_\infty \deq \max_{0 \le i \le t}|x_i|$ is the $\ell^\infty$ norm on $\Reals^{t+1}$.
\end{assumption}

The following qualitative universal approximation result was obtained by \cite{sandberg91structure}: if a causal and time-invariant i/o map $\sF$ satisfies the above two assumptions, then, for any $\eps > 0$, there exists an affine map $A : \Reals^{m+1} \to \Reals^d$ and a lattice map $\ell : \Reals^d \to \Reals$, such that
\begin{align}\label{eq:sandberg_apx}
\sup_{\bdu \in \cM(R)} \sup_{t \in \bbZ_+} \left| (\sF \bdu)_t - \ell \circ A (\bdu_{t-m:t})\right| < \eps,
\end{align}
where we say that a map $\ell : \Reals^d \to \Reals$ is a \textit{lattice map} if $\ell(x_0,\ldots,x_{d-1})$ is generated from $x = (x_0,\ldots,x_{d-1})$ by a finite number of min and max operations that do not depend on $x$. Any lattice map can be implemented using ReLU units, so \eqref{eq:sandberg_apx} is a ReLU TCN approximation guarantee. Our main result is a quantitative version of Sandberg's theorem:

\begin{theorem}\label{thm:main_apx} Let $\sF$ be a causal and time-invariant i/o map satisfying Assumptions~\ref{as:AFM} and \ref{as:cont}.	Then, for any $\eps > 0$ and any $\gamma \in (0,1)$, there exists a ReLU TCN $\wh{\sF}$ with context length $m = m^*_{\sF}(\gamma\eps)$, width $m+2$, and depth $\big( \frac{O(R)}{\omega^{-1}_{m,\sF}((1-\gamma)\eps)}\big)^{m+2}$, such that
	\begin{align}\label{eq:TCN_apx}
		\sup_{\bdu \in \cM(R)} \| \sF\bdu - \wh{\sF}\bdu \|_\infty < \eps.
	\end{align}
\end{theorem}

\begin{remark} {\em The role of the additional parameter $\gamma \in (0,1)$ is to trade off the context length and the depth of the ReLU TCN.}
\end{remark}
\begin{remark} {\em While the approximating ReLU TCN $\wh{\sF}$ is clearly causal, it may not be time-invariant unless $\wh{f}(0,\ldots,0)=0$, where $\wh{f}$ is the ReLU net constructed in the proof below.}
\end{remark}

\begin{proof} Let $m = m^*_\sF(\gamma\eps)$. Since $\tilde{\sF}_m : \Reals^{m+1} \to \Reals$ is continuous with modulus of continuity $\omega_{m,\sF}(\cdot)$, there exists a ReLU net $\wh{f} : \Reals^{m+1} \to \Reals$ of width $m+2$ and depth $\big(\frac{O(R)}{\omega^{-1}_{m,\sF}((1-\gamma)\eps)}\big)^{m+2}$, such that
	\begin{align*}
		\sup_{\bdx \in [-R,R]^{m+1}} |\tilde{\sF}_m(\bdx) - \wh{f}(\bdx)| < (1-\gamma)\eps
	\end{align*}
	\citep{hanin18minwidth}. Consider the TCN $\wh{\sF}$ defined by $(\sF\bdu)_t \deq \wh{f}(u_{t-m},\ldots,u_t)$. Fix an input $\bdu \in \cM(R)$ and consider two cases:

1) If $t \ge m$, then $\bdu_{t-m:t} =  (\sL^{t-m}\sW_{t,m}\bdu)_{0:m}$, where $\sL : \cS \to \cS$ is the left shift operator $(\sL \bdu)_t \deq u_{t+1}$. Therefore,
		\begin{align*}
			(\sF \sW_{t,m}\bdu)_t \stackrel{{\rm (a)}}{=} (\sF \sR^{t-m} \sL^{t-m} \sW_{t,m} \bdu)_t  \stackrel{{\rm (b)}}{=} (\sF \sL^{t-m} \sW_{t,m}\bdu)_m \stackrel{{\rm (c)}}{=} \tilde{\sF}_m(\bdu_{t-m:t}),
		\end{align*}
		where (a) uses the fact that $t \ge m$, (b) is by time invariance of $\sF$, and (c) is by the definition of $\tilde{\sF}_m$.

2) If $t < m$, then $\bdu_{t-m:t} = (\sR^{m-t}\sW_{t,m}\bdu)_{0:m}$ (recall the convention that, for any $\bdv$, we set $v_s = 0$ whenever $s < 0$). Therefore
\begin{align*}
	(\sF \sW_{t,m}\bdu)_t \stackrel{{\rm (a)}}{=} (\sR^{m-t}\sF\sW_{t,m}\bdu)_m \stackrel{{\rm (b)}}{=} (\sF\sR^{m-t}\sW_{t,m}\bdu)_m \stackrel{{\rm (c)}}{=} \tilde{\sF}_m(\bdu_{t-m:t}),
\end{align*}
where (a) uses the fact that $m > t$, (b) is by time invariance, and (c) is by the definition of $\tilde{\sF}_m$.

In either case, the triangle inequality gives
\begin{align*}
	|(\sF\bdu)_t - (\wh{\sF}\bdu)_t| &\le |(\sF\bdu)_t - (\sF\sW_{t,m}\bdu)_t| + |(\sF\sW_{t,m}\bdu)_t - (\wh{\sF}\bdu)_t| \\
	&= |(\sF\bdu)_t - (\sF\sW_{t,m}\bdu)_t| + |\tilde{\sF}_m(\bdu_{t-m:t}) - \wh{f}(\bdu_{t-m:t})| \\
	&< \gamma\eps + (1-\gamma)\eps = \eps.
\end{align*}
Since this holds for all $t$ and all $\bdu$ with $\|\bdu\|_\infty \le R$, the result follows.
\end{proof}

\subsection{The fading memory property}

In order to apply Theorem~\ref{thm:main_apx}, we need control on the context length $m^*_\sF(\cdot)$ and on the modulus of continuity $\omega_{t,\sF}(\cdot)$. In general, these quantities are difficult to estimate. However, it was shown by \citet{park1992criteria} that the property of approximately finite memory is closely related to the notion of \textit{fading memory}, first introduced by \citet{boyd85fading}. Intuitively, an i/o map $\sF$ has fading memory if the outputs at any time $t$ due to any two inputs $\bdu$ and $\bdv$ that were close to one another in recent past will also be close.

Let $\cW$ denote the subset of $\cS$ consisting of all sequences $\bd{w}$, such that $w_t \in (0,1]$ for all $t$ and $w_t \downarrow 0$ as $t \to \infty$. We will refer to the elements of $\cW$ as \textit{weighting sequences}. Then we have the following definition, due to \cite{park1992criteria}:

\begin{definition} We say that an i/o map $\sF$ has {\em fading memory} on $\cM \subseteq \cS$ with respect to $\bd{w} \in \cW$ if for any $\eps > 0$ there exists $\delta > 0$ such that, for all $\bdu,\bdv \in \cM$ and all $t \in \bbZ_+$,
	\begin{align}\label{eq:fading}
		\max_{s \in \{0,\ldots,t\}} w_{t-s}|u_s - v_s| < \delta \quad \Longrightarrow \quad |(\sF\bdu)_t-(\sF\bdv)_t| < \eps.
	\end{align}
\end{definition}
The weighting sequence $\bdw$ governs the rate at which the past values of the input are discounted in determining the current output. To capture the best trade-offs in \eqref{eq:fading}, we will also use a $\bdw$-dependent modulus of continuity:
\begin{align*}
	\alpha_{\bdw,\sF}(\delta) \deq \sup \Big\{ |(\sF\bdu)_t - (\sF\bdv)_t| : t \in \bbZ_+, \bdu,\bdv \in \cM, \max_{s \in \{0,\ldots,t\}} w_{t-s}|u_s-v_s| \le \delta\Big\}.
\end{align*}
It was shown by \citet{park1992criteria} that an i/o map satisfies Assumptions~\ref{as:AFM} and \eqref{as:cont} if and only if it has fading memory with respect to some (and hence any) $\bdw \in \cW$. The following result provides a quantitative version of this equivalence:
\begin{proposition}\label{prop:fading_afm} Let $\sF$ be an i/o map.
	\begin{enumerate}
		\item If $\sF$ satisfies Assumptions~\ref{as:AFM} and \ref{as:cont}, then it has fading memory on $\cM$ with respect to any weighting sequence $\bdw \in \cW$, and
		\begin{align}\label{eq:AFM_to_w}
			\alpha^{-1}_{\bdw,\sF}(\eps) \ge w_{m^*_\sF(\eps/3)} \omega^{-1}_{m^*_\sF(\eps/3),\sF}(\eps/3).
		\end{align}
		\item If $\sF$ has fading memory on $\cM(R)$ with respect to some $\bdw \in \cW$, then it has satisfies Assumptions~\ref{as:AFM} and \ref{as:cont}, and
		\begin{align}\label{eq:w_to_AFM}
			m^*_\sF(\eps; R) \le  \inf \Big\{ m \in \bbZ_+ : w_m \le  \frac{\alpha^{-1}_{\bdw,\sF}(\eps)}{R}\Big\} \qquad \text{and} \qquad \omega_{t,\sF}(\delta) \le \alpha_{\bdw,\sF}(\delta).
		\end{align}
	\end{enumerate}

\end{proposition}

\section{Recurrent systems}
\label{sec:recurrent}

So far, we have considered arbitrary i/o maps $\sF : \cS \to \cS$. However, many such maps admit \textit{state-space realizations} \citep{sontag98control} --- there exist a state transition map $f : \Reals^n \times \Reals \to \Reals^n$, an output map $g : \Reals^n \to \Reals$, and an initial condition $\xi \in \Reals^n$, such that the output sequence $\bdy = \sF \bdu$ is detemined recursively by
\begin{subequations}\label{eq:recurrent}
	\begin{align}
		x_{t+1} &= f(x_t,u_t) \\
		y_t &= g(x_t)
	\end{align}
\end{subequations}
with $x_0 = \xi$. The i/o map $\sF$ realized in this way is evidently causal, and it is time-invariant if $f(\xi,0) = \xi$ and $g(\xi) = 0$. In this section, we will identify the conditions under which recurrent models satisfy Assumptions~\ref{as:AFM} and \ref{as:cont}. Along the way, we will derive the approximation results of \citet{miller2019stable} as a special case.

\subsection{Approximately finite memory and incremental stability}

Consider the system in \eqref{eq:recurrent}. Given any input $\bdu \in \cS$, any $\xi \in \Reals^n$, and any $s,t \in \bbZ_+$ with $t \ge s$, we denote by $\varphi^{\bdu}_{s,t}(\xi)$ the state at time $t$  when $x_s = \xi$.  Let $\cM$ be a subset of $\cS$.  We say that $\bbX \subseteq \Reals^n$ is a \textit{positively invariant set} of \eqref{eq:recurrent} for inputs in $\cM$ if, for all $\xi \in \bbX$, all $\bdu \in \cM$, and all $0 \le s \le t$, $\varphi^{\bdu}_{s,t}(\xi) \in \bbX$. We will be interested in systems with the following property \citep{tran2017convergent}:

\begin{definition} The system \eqref{eq:recurrent} is {\em uniformly asymptotically incrementally stable} for inputs in $\cM$ on a positively invariant set $\bbX$ if there exists a function $\beta : \Reals_+ \times \Reals_+ \to \Reals_+$ of class $\cK\cL$\footnote{A function $\beta : \Reals_+ \times \Reals_+ \to \Reals_+$ is of class $\cK\cL$ if it is continuous and strictly increasing in its first argument, continuous and strictly decreasing in its second argument, $\beta(0,t)=0$ for any $t$, and $\lim_{t \to \infty}\beta(r,t) = 0$ for any $r$ \citep{sontag98control}.}, such that the inequality
\begin{align}\label{eq:UAIS}
	\| \varphi^{\bdu}_{s,t}(\xi) - \varphi^{\bdu}_{s,t}(\xi')\| \le \beta( \|\xi - \xi'\|, t-s)
\end{align}
holds for all inputs $\bdu \in \cM$, all initial conditions $\xi,\xi' \in \bbX$, and all $0 \le s \le t$, where $\|\cdot\|$ is the $\ell^2$ norm on $\Reals^n$.
\end{definition}
In other words, a system is incrementally stable if the influence of any initial condition in $\bbX$ on the state trajectory is asymptotically negligible. A key consequence  is the following estimate:
	\begin{proposition}\label{prop:UAIS_error_est} Let $\bdu,\tilde{\bdu}$ be two input sequences in $\cM$. Then, for any $\xi \in \bbX$ and any $t \in \bbZ_+$,
		\begin{align}\label{eq:UAIS_error_est}
			\| \varphi^{\bdu}_{0,t}(\xi) - \varphi^{\tilde{\bdu}}_{0,t}(\xi) \| \le \sum^{t-1}_{s=0} \beta\left( \| f(\tilde{x}_s,u_s) - f(\tilde{x}_s,\tilde{u}_s)\|, t-s-1\right),
		\end{align}
		where $x_s$ and $\tilde{x}_s$ denote the states at time $s$ due to inputs $\bdu$ and $\tilde{\bdu}$, respectively, with $x_0 = \tilde{x}_0 = \xi$.
	\end{proposition}
	
Consider a state-space model \eqref{eq:recurrent} with a positively invariant set $\bbX$, with the following assumptions:

\begin{assumption}\label{as:Lipschitz} The state transition map $f(x,u)$ is $L_f$-Lipschitz in $u$ for all $x \in \bbX$ and the output map $g(x)$ is $L_g$-Lipschitz in $x \in \bbX$.
\end{assumption}

\begin{assumption}\label{as:compact} For any initial condition $\xi \in \bbX$ there exists a compact set $\bbS_\xi \subseteq \bbX$ such that $\varphi^{\bdu}_{0,t}(\xi) \in \bbS_\xi$ for all $\bdu \in \cM(R)$ and all $t \in \bbZ_+$.
\end{assumption}

\begin{assumption}\label{as:UAIS} The system \eqref{eq:recurrent} is uniformly asymptotically incrementally stable on $\bbX$ for inputs in $\cM(R)$, and the function $\beta$ in \eqref{eq:UAIS} satisfies the summability condition
	\begin{align}\label{eq:summability}
		\sum_{t \in \bbZ_+} \beta(C,t) < \infty
	\end{align}
for any $C \ge 0$. (For example, if $\beta(C,k) = Ck^{-\alpha}$ for some $\alpha > 1$, then this condition is satisfied.)
\end{assumption}

We are now in position to prove the main result of this section:

\begin{theorem}\label{thm:UAIS_AFM} Suppose that Assumptions~\ref{as:Lipschitz}--\ref{as:UAIS} are satisfied. Then the i/o map $\sF$ of the system \eqref{eq:recurrent} satisfies Assumptions~\ref{as:AFM} and \ref{as:cont} with
	\begin{align}\label{eq:meps_UAIS}
		m^*_\sF(\eps) \le \min \Big\{ m \in \bbZ_+ : \sum_{k \ge m} \beta({\rm diam}(\bbS_\xi),k) < \eps/L_g \Big\}
	\end{align}
	and
	\begin{align}\label{eq:cont_UAIS}
		\omega_{t,\sF}(\delta) \le L_g\sum^{t-1}_{s=0} \beta (L_f\delta,s), \qquad \forall t \in \bbZ_+.
	\end{align}
\end{theorem}

\begin{proof} Fix some $t,m \in \bbZ_+$. For an arbitrary input $\bdu \in \cM(R)$, let $\tilde{\bdu} = \sW_{t,m}\bdu$, where we may assume without loss of generality that $t \ge m$. Then $\tilde{u}_s = u_s \1_{\{t-m \le s \le t\}}$, and therefore
\begin{align}
	 \sum^{t-1}_{s=0} \beta\left( \| f(\tilde{x}_s,u_s) - f(\tilde{x}_s,\tilde{u}_s)\|, t-s-1\right) &= \sum^{t-m-1}_{s = 0}  \beta\left( \| f(\tilde{x}_s,u_s) - f(\tilde{x}_s,0)\|, t-s-1\right) \nonumber\\
	 &\le \sum^{t-m-1}_{s=0} \beta({\rm diam}(\bbS_\xi), t - s - 1) \nonumber\\
	 &\le \sum^\infty_{s=m} \beta({\rm diam}(\bbS_\xi),s).\label{eq:betasum_UAIS}
\end{align}
By the summability condition \eqref{eq:summability}, the summation in \eqref{eq:betasum_UAIS} converges to $0$ as $m \uparrow \infty$. Thus, if we choose $m$ so that the right-hand side of \eqref{eq:betasum_UAIS} is smaller than $\eps/L_g$, it follows from Proposition~\ref{prop:UAIS_error_est}  that
\begin{align*}
	|(\sF\bdu)_t-(\sF\sW_{t,m}\bdu)_t| &= |g(\varphi^{\bdu}_{0,t}(\xi))-g(\varphi^{\tilde{\bdu}}_{0,t}(\xi))| \le L_g \| \varphi^{\bdu}_{0,t}(\xi) - \varphi^{\tilde{\bdu}}_{0,t}(\xi)| < \eps.
\end{align*}
This proves \eqref{eq:meps_UAIS}. Now fix any two $\bdu,\tilde{\bdu}\in \cM(R)$ with $\|\bdu_{0:t}-\tilde{\bdu}_{0:t}\|_\infty < \delta$. Then $\max_{0 \le s \le t}\|f(x,u_s)-f(x,\tilde{u}_s)\| \le L_f \delta$ for all $x \in \bbX$, so Proposition~\ref{prop:UAIS_error_est} gives
\begin{align*}
	|\tilde{\sF}_t(\bdu_{0:t})-\tilde{\sF}_t(\tilde{\bdu}_{0:t})| &= |g(\varphi^{\bdu}_{0,t}(\xi))-g(\varphi^{\tilde{\bdu}}_{0,t}(\xi))| \\
	&\le L_g \| \varphi^{\bdu}_{0,t}(\xi) - \varphi^{\tilde{\bdu}}_{0,t}(\xi)\| \\
	&\le L_g \sum^{t-1}_{s=0} \beta(L_f \delta,s),
\end{align*}
which proves \eqref{eq:cont_UAIS}. \end{proof}

\subsection{Exponential incremental stability and the Demidovich criterion}

\citet{miller2019stable} consider the case of contracting systems: there exists some $\lambda \in (0,1)$ and a set $\bbU \subseteq \Reals^m$, such that
\begin{align}\label{eq:contractive}
	\|f(x,u)-f(x',u)\| \le \lambda\|x-x'\|
\end{align}
for all $x,x' \in \Reals^n$ and all $u \in \bbU$. Such a system is \textit{uniformly exponentially incrementally stable} on any positively invariant set $\bbX$, with $\beta(C,t) = C\lambda^t$. In this section, we obtain their result as a special case of a more general stability criterion, known in the literature on nonlinear system stability as the \textit{Demidovich criterion} \citep{pavlov06nonlinea}. The following result is a simplified version of a more general result of \citet{tran2017convergent}:

\begin{proposition}[the discrete-time Demidovich criterion]\label{prop:Demidovich} Consider the recurrent system \eqref{eq:recurrent} with a convex positively invariant set $\bbX$, where the state transition map $f(x,u)$ is differentiable in $x$ for any $u \in \bbU$. Suppose that there exists a symmetric positive definite matrix $P$ and a constant $\mu \in (0,1)$, such that
	\begin{align}\label{eq:Demidovich}
		\frac{\partial}{\partial x}f(x,u)^\top P \frac{\partial}{\partial x} f(x,u) - \mu P \preceq 0
	\end{align}
	for all $x \in \bbX$ and all $u \in \bbU$, where $\frac{\partial}{\partial x}f(x,u)$ is the Jacobian of $f(\cdot,u)$ with respect to $x$. Then the system \eqref{eq:recurrent} is uniformly exponentially incrementally stable with $\beta(C,t) = \sqrt{\kappa(P)} C \mu^{t/2}$, where $\kappa(P)$ is the condition number of $P$. 
\end{proposition}

\begin{proof} Fix any $u \in \bbU$ and $\xi,\xi' \in \bbX$, and define the function $\Phi : [0,1] \to \Reals$ by
	\begin{align*}
		\Phi(s) \deq (f(\xi,u)-f(\xi',u))^\top P f(s \xi + (1-s)\xi',u).
	\end{align*}
Then 
\begin{align}\label{eq:Phidiff_1}
	\Phi(1)-\Phi(0) = (f(\xi,u)-f(\xi',u))^\top P (f(\xi,u)-f(\xi',u)).
\end{align}
By the mean-value theorem, there exists some $\bar{s} \in [0,1]$, such that
\begin{align}\label{eq:Phidiff_2}
	\Phi(1)-\Phi(0) = \frac{\dif}{\dif s} \Phi(s)\Big|_{s = \bar{s}} = (f(\xi,u)-f(\xi',u))^\top P \frac{\partial}{\partial x}f(\bar{\xi},u)(\xi-\xi'),
\end{align}
where $\bar{\xi} = \bar{s}\xi + (1-\bar{s})\xi' \in \bbX$, since $\bbX$ is convex. From \eqref{eq:Demidovich}, \eqref{eq:Phidiff_1}, and \eqref{eq:Phidiff_2} it follows that
\begin{align*}
&	(f(\xi,u)-f(\xi',u))^\top P (f(\xi,u)-f(\xi',u)) \nonumber\\
&\qquad\le (\xi-\xi')^\top \frac{\partial}{\partial x}f(\bar{\xi},u)^\top P \frac{\partial}{\partial x}f(\bar{\xi},u)(\xi-\xi') \\
&\qquad\le \mu (\xi-\xi')^\top P (\xi-\xi').
\end{align*}
Define the function $V : \bbX \times \bbX \to \Reals_+$ by $V(\xi,\xi') \deq (\xi-\xi')^\top P (\xi-\xi')$. From the above estimate, it follows that $V$ is a \textit{Lyapunov function} for the dynamics, i.e., for any $u \in \bbU$ and $\xi,\xi' \in \bbX$,
\begin{align}\label{eq:Lyapunov}
	V(f(\xi,u),f(\xi',u)) \le \mu V(\xi,\xi').
\end{align}
Consequently, for any input $\bdu$ with $u_t \in \bbU$ for all $t$ and any $\xi,\xi' \in \bbX$,
\begin{align*}
	V(\varphi^{\bdu}_{0,t+1}(\xi),\varphi^{\bdu}_{0,t+1}(\xi')) &= V(f(\varphi^{\bdu}_{0,t}(\xi),u_t),f(\varphi^{\bdu}_{0,t}(\xi'),u_t)) \\
	&\le \mu V(\varphi^{\bdu}_{0,t}(\xi),\varphi^{\bdu}_{0,t}(\xi')).
\end{align*}
Iterating, we obtain the inequality $V(\varphi^{\bdu}_{0,t}(\xi),\varphi^{\bdu}_{0,t}(\xi')) \le \mu^t V(\xi,\xi')$. Finally, since $P \succ 0$,
\begin{align*}
	\|\varphi^{\bdu}_{0,t}(\xi)-\varphi^{\bdu}_{0,t}(\xi)\|^2 \le \frac{\lambda_{\max}(P)}{\lambda_{\min}(P)}\mu^t \|\xi-\xi'\|^2 = \kappa(P)\|\xi-\xi'\|^2 \mu^t,
\end{align*}
and the proof is complete.
\end{proof}

\begin{theorem}\label{thm:recurrent} Suppose the system \eqref{eq:recurrent} satisfies Assumption~\ref{as:Lipschitz} and the Demidovich criterion with $\bbU = [-R,R]$, its positively invariant set $\bbX$ contains $0$, and $f(0,0) = 0$. Then its i/o map $\sF$ with zero initial condition $x_0 = 0$ satisfies Assumptions~\ref{as:AFM} and \ref{as:cont} with
\begin{align}\label{eq:exp_stable_AFM}
	m^*_\sF(\eps) \le \frac{2\log (\frac{2\kappa(P)L_fL_gR}{(1-\sqrt{\mu})^2\eps})}{\log \frac{1}{\mu}} \qquad \text{and} \qquad \omega_{t,\sF}(\delta) \le \frac{\sqrt{\kappa(P)}L_fL_g\delta}{1-\sqrt{\mu}}.
\end{align}
\end{theorem}
\begin{proof} Since $P$ is symmetric and positive definite, $\|x\|_P \deq \sqrt{x^\top P x}$ is a norm on $\Reals^n$ with $\lambda_{\min}(P)\|\cdot\|^2 \le \|\cdot\|^2_P \le \lambda_{\max}(P)\|\cdot\|^2$. Then, for all $\xi \in \bbX$, $\bdu \in \cM(R)$, and $t$,
	\begin{align*}
		\|\varphi^{\bdu}_{0,t+1}(\xi)\|_P &= \|f(\varphi^{\bdu}_{0,t}(\xi),u_t) \|_P \\
		&\le \|f(\varphi^{\bdu}_{0,t}(\xi),u_t)-f(0,u_t) \|_P + \|f(0,u_t) - f(0,0)\|_P \\
		&\le \sqrt{\mu} \|\varphi^{\bdu}_{0,t}(\xi) \|_P + \sqrt{\lambda_{\max}(P)} L_f R,
	\end{align*}
	where we have used the Lyapunov bound \eqref{eq:Lyapunov}. Unrolling the recursion gives the estimate
	\begin{align*}
		\sup_{t \in \bbZ_+} \sup_{\bdu \in \cM(R)}\|\varphi^{\bdu}_{0,t}(\xi)\|_P \le \sqrt{\mu}\|\xi\|_P + \frac{\sqrt{\lambda_{\max}(P)}L_fR}{1-\sqrt{\mu}}.
	\end{align*}
Thus, Assumption~\ref{as:compact} is satisfied, where $\bbS_\xi$ is the ball of $\ell^2$-radius $\sqrt{\kappa(P)}\left(\|\xi\| + \frac{L_fR}{1-\sqrt{\mu}}\right)$ centered at $0$. Assumption~\ref{as:UAIS} is also satisfied by Proposition~\ref{prop:Demidovich}. The estimates in \eqref{eq:exp_stable_AFM} follow from Theorem~\ref{thm:UAIS_AFM}.
\end{proof}
The following result now follows as a direct consequence of Theorems~\ref{thm:main_apx} and \ref{thm:recurrent}:
\begin{corollary} If the system \eqref{eq:recurrent} satisfies the conditions of Theorem~\ref{thm:recurrent}, then its i/o map $\sF$ with zero initial condition can be $\eps$-approximated in the sense of Theorem~\ref{thm:main_apx} by a ReLU TCN $\wh{\sF}$ with width ${\rm polylog}(\frac{1}{\eps})$ and depth ${\rm quasipoly}(\frac{1}{\eps})$.\footnote{We say that a given quantity $N$ has \textit{quasipolynomial growth} in $1/\eps$, and write $N \le {\rm quasipoly}(1/\eps)$, if $N = O(\exp({\rm polylog}(\frac{1}{\eps})))$.}
\end{corollary}

\subsection{Contractivity vs.\ the Demidovich criterion}

If the contractivity condition \eqref{eq:contractive} holds and $f(x,u)$ is differentiable in $x$, then the Demidovich criterion is satisfied with $P = I_n$ and $\mu =\lambda^2$. In that case, we immediately obtain the exponential estimate $\beta(C,t) \le C\lambda^t$. However, the Demidovich criterion covers a wider class of nonlinear systems. As an example, consider a discrete-time nonlinear system of \textit{Lur'e type} (cf.~\citet{sandberg93circle,kim14lure,sarkans2016lure} and references therein):
	\begin{subequations}\label{eq:Lur'e}
	\begin{align}
		x_{t+1} &= Ax_t + B\psi(u_t - y_t) \\
		y_t &= Cx_t
	\end{align}
	\end{subequations}
	Here, the state $x_t$ is $n$-dimensional while the input $u_t$ and the output $y_t$ are scalar, so $A \in \Reals^{n \times n}$, $B \in \Reals^{n \times 1}$, and $C \in \Reals^{1 \times n}$. The map $\psi : \Reals \to \Reals$ is a fixed differentiable nonlinearity. The system in \eqref{eq:Lur'e} has the form \eqref{eq:recurrent} with $f(x,u) = Ax + B\psi(u-Cx)$ and $g(x) = Cx$, and can be realized as the negative feedback interconnection of the discrete-time linear system
	\begin{subequations}\label{eq:LTI}\begin{align}
		x_{t+1} &= Ax_t + Bv_t \\
		y_t &= Cx_t
	\end{align}
	\end{subequations}
and the nonlinear element $\psi$ using the feedback law $v_t = \psi(u_t - y_t)$. We make the following assumptions (see, e.g., \citet{sontag98control} for the requisite control-theoretic background):
	\begin{assumption}\label{as:Lur'e1} The nonlinearity $\psi : \Reals \to \Reals$ satisfies $\psi(0)=0$, and there exist real numbers $-\infty < a \le b < \infty$ such that $a \le \psi'(\cdot) \le b$.
	\end{assumption}
	\begin{assumption}\label{as:Lur'e2} $A$ is a {\em Schur matrix}, i.e., its spectral radius $\rho(A)$ is strictly smaller than $1$; the pair $(A,B)$ is {\em controllable}, i.e., the $n\times n$ matrix $[  B \,|\, AB \,|\, \ldots \,|\, A^{n-1}B]$ has rank $n$; and the pair $(A,C)$ is {\em observable}, i.e., the $n \times n$ matrix $[  C^\top \,|\, A^\top C^\top \,|\, \dots  \,|\, (A^\top)^{n-1}C^\top ]$ has rank $n$. 
	\end{assumption}
	\begin{assumption}\label{as:Lur'e3} Let ${\mathbb T} \deq \{z \in {\mathbb C} : |z|=1\}$ denote the unit circle in the complex plane. The rational function $G(z) \deq C(zI_n-A)^{-1}B$  satisfies
		\begin{align}\label{eq:Hinfty}
			\|G\|_{\cH_\infty({\mathbb T})} \deq \sup_{z \in {\mathbb T}} |G(z)| < \gamma^{-1}
		\end{align}
		for some $\gamma > 0$ such that $r^2 \le \gamma^2$ for all $a \le r \le b$.
	\end{assumption}
\begin{remark}{\em Assumption~\ref{as:Lur'e1} imposes a {\em slope condition} on $\psi$ and is standard in the analysis of Lur'e systems \citep{tsypkin64circle,sandberg91structure,kim14lure}. The function $G(z)$ is the {\em transfer function} of the linear system \eqref{eq:LTI}. Assumption~\ref{as:Lur'e2} states that the triple $(A,B,C)$ is a {\em minimal realization} of $G$. The quantity $\|G\|_{\cH_\infty({\mathbb T})}$ appearing in Eq.~\eqref{eq:Hinfty} in Assumption~\ref{as:Lur'e3} is the {\em $\cH_\infty$-norm} of $G$ on the unit circle  in the complex plane. Assumptions~\ref{as:Lur'e2} and \ref{as:Lur'e3} are also common and are in the spirit of the well-known {\em circle criterion} \citep{tsypkin64circle,sandberg93circle}.}
\end{remark}
With these preliminaries out of the way, we have the following:
\begin{proposition}\label{prop:Lur'e} Suppose that the system \eqref{eq:Lur'e} satisfies Assumptions~\ref{as:Lur'e1}--\ref{as:Lur'e3}. Then it satisfies the discrete-time Demidovich criterion with $\bbX = \Reals^n$ and $\bbU = \Reals$, and moreover $\mu > \rho(A)^2$.
\end{proposition}
The crucial ingredient in the proof is the Discrete-Time Bounded-Real Lemma \citep{vaidyanathan85dtbr}, which guarantees the existence of the matrix $P$ appearing in the Demidovich criterion. The main takeaway here is that the function $f(x,u) = Ax+B\psi(u-Cx)$ need not be contractive (i.e., it may be the case that $P \neq I_n$), but it will be contractive in the $\|\cdot\|_P$ norm.

\section{Comparison of architectures}
\label{sec:comparison}

So far, we have shown that any i/o map $\sF$ with approximately finite memory can be approximated by a ReLU temporal convolutional net. We have also considered recurrent models and shown that any incrementally stable recurrent model has approximately finite memory and can therefore be approximated by a ReLU TCN. As far as their approximation capabilities are concerned, both recurrent models and autoregressive models like TCNs are equivalent, since any finite-memory i/o map of the form \eqref{eq:FM} admits the state-space realization
\begin{align*}
	& x^1_{t+1} = x^2_t, 
	x^2_{t+1} = x^3_t, \ldots, 
	x^{m-1}_{t+1} = x^m_t,
	x^m_{t+1} = u_t \\
	& y_t = f(x^1_t,x^2_t,\ldots,x^m_t,u_t)
\end{align*}
of the tapped delay line type, with zero initial condition $(x^1_0,\ldots,x^m_0) = (0,\ldots,0)$. (Compared to \eqref{eq:recurrent}, we are allowing a direct `feedthrough' connection from the input $u_t$ to the output $y_t$.) The advantage of autoregressive models like TCNs shows up during training and regular operation, since shifted copies of the input sequence can be efficiently processed in parallel rather than sequentially.

Another point worth mentioning is that, while the construction in the proof of Theorem~\ref{thm:main_apx} makes use of ReLU nets as a universal function approximator, any other family of universal approximators can be used instead, for example, multivariate polynomials or rational functions. In fact, if one uses multivariate polynomials to approximate the functionals $\tilde{\sF}_t$, the resulting family of i/o maps is known as the (discrete-time) finite Volterra series \citep{boyd85fading}, and has been used widely in the analysis of nonlinear systems. However, TCNs generally provide a more parsimonious representation. To see this, consider the following (admittedly contrived) example of an i/o map:
\begin{align}\label{eq:ReLU_example}
	(\sF\bdu)_t = \ReLU\Bigg(\sum^\infty_{s=0} h_s u_{t-s}\Bigg),
\end{align}
where the filter coefficients $h_t$ have the exponential decay property $|h_t| \le C\lambda^t$ for some $C > 0$ and $\lambda \in (0,1)$. It is not hard to show that $\sF$ has exponentially fading memory, and a very simple $\eps$-approximation by a TCN is obtained by zeroing out all of the filter coefficients $h_s,s > m \sim \log (\frac{1}{\eps})$:
\begin{align*}
	(\wh{\sF}\bdu)_t = \ReLU\Bigg(\sum^m_{s=0}h_s u_{t-s}\Bigg).
\end{align*}
However, any $\eps$-approximation for $\sF$ using Volterra series would need ${\rm poly}(\frac{1}{\eps})$ terms, since the best polynomial $\eps$-approximation of the ReLU on any compact interval has degree $\Omega(\frac{1}{\eps})$ \citep[Chap.~9, Thm.~3.3]{devore1993apx}. On the other hand, if we consider an i/o map of the form \eqref{eq:ReLU_example}, but with a degree-$d$ univariate polynomial instead of ReLU, then we can $\eps$-approximate it with a TCN of depth $O(d + \log \frac{d}{\eps})$ and $O(d \log \frac{d}{\eps})$ units \citep{liang2017deep}.

\subsubsection*{Acknowledgments} This work was supported in part by the National Science Foundation under the Center for Advanced Electronics through Machine Learning (CAEML) I/UCRC award no.~CNS-16-24811. 


\bibliography{afm.bib,references.bib}

\newpage

\appendix
\section{Omitted proofs}

\setcounter{equation}{0}
\renewcommand\theequation{A.\arabic{equation}}

\begin{proof}[Proof of Proposition~\ref{prop:fading_afm}] Suppose $\sF$ satisfies Assumptions~\ref{as:AFM} and \ref{as:cont}. Fix some $\eps > 0$ and let $m = m^*_\sF(\eps/3)$ and $\delta = w_m\omega^{-1}_{m,\sF}(\eps/3)$. Now fix some $t \in \bbZ_+$ and consider any two $\bdu,\bdv \in \cM(R)$ such that
	\begin{align}\label{eq:dist_uv}
	\max_{s \in \{0,\ldots,t\}} w_{t-s}|u_s - v_s| < \delta.
	\end{align}
Using the same reasoning as in the proof of Theorem~\ref{thm:main_apx}, we can write $(\sF\sW_{t,m}\bdu)_t = \tilde{\sF}_m(\bdu_{t-m:t})$ and $(\sF\sW_{t,m}\bdv)_t = \tilde{\sF}_m(\bdv_{t-m:t})$, where, as before, we set $u_s = v_s = 0$ for $s < 0$. From the monotonicity of $\bdw$ and \eqref{eq:dist_uv} it follows that
\begin{align*}
	\| \bdu_{t-m:t} - \bdv_{t-m:t} \|_\infty \le \frac{1}{w_m}\max_{s \in \{t-m,\ldots,t\}} w_{t-s}|u_s - v_s| < \omega^{-1}_{m,\sF}(\eps/3),
\end{align*}
which implies that
\begin{align*}
	|(\sF \sW_{t,m}\bdu)_t - (\sF \sW_{t,m}\bdv)_t| &= |\tilde{\sF}_m(\bdu_{t-m:t})-\tilde{\sF}_m(\bdv_{t-m:t})| < \eps/3.
\end{align*}
Altogether, we see that \eqref{eq:dist_uv} implies that
\begin{align*}
	|(\sF \bdu)_t - (\sF \bdv)_t| &\le |(\sF\bdu)_t - (\sF\sW_{t,m}\bdu)_t| + |(\sF \sW_{t,m}\bdu)_t - (\sF \sW_{t,m}\bdv)_t| + |(\sF\bdv)_t - (\sF\sW_{t,m}\bdv)_t| \nonumber\\
	&< \eps/3 + \eps/3 + \eps/3 = \eps,
\end{align*}
which leads to \eqref{eq:AFM_to_w}.

Now suppose that $\sF$ has fading memory w.r.t.\ $\bdw$. Given $\eps > 0$, let $\delta = \alpha^{-1}_{\bdw,\sF}(\eps)$ and choose any $m \in \bbZ_+$, such that $w_m < \delta/R$. If $t < m$, then $\bd{u}_{0:t} = (\sW_{t,m}\bdu)_{0:t}$, and thus $(\sF\bdu)_t = (\sF\sW_{t,m}\bdu)_t$. On the other hand, if $t \ge m$, then, for any $\bdu \in \cM(R)$,
	\begin{align*}
		\max_{s \in \{0,\ldots,t\}} |u_s - (\sW_{t,m} \bdu)_s| = \begin{cases}
		0, & t-m \le s \le t \\
		|u_s|, & s < t-m
		\end{cases}
	\end{align*}
	and therefore, by the monotonicity of $\bdw$ and the choice of $m$,
	\begin{align*}
		\max_{s \in \{0,\ldots,t\}} w_{t-s} |u_s - (\sW \bdu_{t,m})_s| &= \max_{s < t-m} w_{t-s}|u_s| \le w_m \|\bdu\|_\infty < \delta,
	\end{align*}
	which implies that $|(\sF\bdu)_t - (\sF\sW_{t,m}\bdu)_t| < \eps$. Consequently, $m^*_\sF(\eps) \le m$. Moreover, since the elements of $\bdw$ take values in $(0,1]$, it follows from definitions that, for any $\bdu,\bdv \in \cM(R)$ and any $t$, 
	\begin{align*}
		\|\bdu_{0:t} - \bdv_{0:t}\|_\infty < \delta \quad \Longrightarrow \quad \max_{s \in \{0,\ldots,t\}} w_{t-s}|u_s-v_s| < \delta \quad \Longrightarrow \quad |(\sF\bdu)_t - (\sF\bdv)_t| \le \alpha_{\bdw,\sF}(\delta).
	\end{align*}
This establishes \eqref{eq:w_to_AFM}.
\end{proof}

	\begin{proof}[Proof of Proposition~\ref{prop:UAIS_error_est}]
The family of mappings $\varphi^{\bdu}_{s,t}(\cdot)$ has the following \textit{semiflow property}: for any input $\bdu$ and any $0 \le r \le s \le t$,
\begin{align}\label{eq:semiflow}
	\varphi^{\bdu}_{r,t}(\xi) = \varphi^{\bdu}_{s,t}(\varphi^{\bdu}_{r,s}(\xi)).
\end{align}		
By telescoping and by the semiflow property \eqref{eq:semiflow}, we have
		\begin{align}
			\varphi^{\bdu}_{0,t}(\xi) - \varphi^{\tilde{\bdu}}_{0,t}(\xi) &= \sum^{t-1}_{s=0} \left(\varphi^{\bdu}_{s,t}(\varphi^{\tilde{\bdu}}_{0,s}(\xi))-\varphi^{\bdu}_{s+1,t}(\varphi^{\tilde{\bdu}}_{0,s+1}(\xi))\right) \nonumber\\
			&= \sum^{t-1}_{s=0} \left(\varphi^{\bdu}_{s+1,t}(\varphi^{\bdu}_{s,s+1}(\varphi^{\tilde{\bdu}}_{0,s}(\xi)))-\varphi^{\bdu}_{s+1,t}(\varphi^{\tilde{\bdu}}_{0,s+1}(\xi))\right).\label{eq:telescoping}
		\end{align}
Using the fact that $\varphi^{\bdu}_{s,s+1}(\varphi^{\tilde{\bdu}}_{0,s}(\xi)) = \varphi^{\bdu}_{s,s+1}(f(\varphi^{\tilde{\bdu}}_{0,s}(\xi),u_s))$ and the stability property \eqref{eq:UAIS}, 
\begin{align*}
	\left\| \varphi^{\bdu}_{s+1,t}(\varphi^{\bdu}_{s,s+1}(\varphi^{\tilde{\bdu}}_{0,s}(\xi)))-\varphi^{\bdu}_{s+1,t}(\varphi^{\tilde{\bdu}}_{0,s+1}(\xi))\right\| \le \beta\left( \| f(\tilde{x}_s,u_s) - f(\tilde{x}_s,\tilde{u}_s)\|, t-s-1\right).
\end{align*}
Substituting this into \eqref{eq:telescoping}, we get \eqref{eq:UAIS_error_est}.
\end{proof}

\begin{proof}[Proof of Proposition~\ref{prop:Lur'e}] Since the matrix $A$ is Schur, the function
	\begin{align*}
		g(r) \deq \sup_{z \in {\mathbb T}} |G(rz)| = \|G(r\cdot)\|_{\cH_\infty({\mathbb T})}, \qquad r > \rho(A)
	\end{align*}
is continuous.  In particular, there exists some $r_0 \in (\rho(A),1)$, such that $g(r_0) < g(1) < \gamma^{-1}$. Consequently, the rational function
\begin{align*}
	H(z) \deq \gamma G(r_0z) = \frac{\gamma C}{r_0}\left(zI_n - \frac{A}{r_0}\right)^{-1} B
\end{align*}
is well-defined for all $z \in {\mathbb C}$ with $|z| \ge r_0$, and we have the following:
\begin{itemize}
	\item $\frac{A}{r_0}$ is a Schur matrix;
	\item the pair $(\frac{A}{r_0},B)$ is controllable;
	\item the pair $(\frac{A}{r_0}, \frac{\gamma C}{r_0})$ is observable;
	\item $\|H\|_{\cH_\infty({\mathbb T})} < 1$.
\end{itemize}
Then, by the Discrete-Time Bounded-Real Lemma \citep{vaidyanathan85dtbr}, there exist real matrices $L,W$ and a symmetric positive definite matrix $P \in \Reals^{n \times n}$, such that
\begin{subequations}\label{eq:DTBR}
	\begin{align}
		A^\top P A + \gamma^2 C^\top C + r^2_0 L^\top L &= r^2_0 P \\
		B^\top P B + W^\top W &= I_n \\
		A^\top P B + r_0 L^\top W &= r_0 I_n.
	\end{align}
\end{subequations}
From \eqref{eq:DTBR}, for any $\theta \in \Reals$ we have
\begin{align*}
&	(A-\theta BC)^\top P (A-\theta BC) - r^2_0 P\nonumber\\
 &\qquad = A^\top P A - \theta(C^\top B^\top P A + A^\top P BC) + \theta^2 C^\top B^\top P B C - r^2_0P \\
 &\qquad = (\theta^2 - \gamma^2) C^\top C - (r_0 L - \theta WC)^\top (r_0 L - \theta WC).
\end{align*}
Let $\mu \deq r^2_0$. Then, since $\gamma^2 \ge \theta^2$ for all $\theta \in [a,b]$, it follows that
\begin{align*}
	(A - \theta BC)^\top P (A-\theta BC) - \mu P \preceq 0, \qquad a \le \theta \le b.
\end{align*}
Since
\begin{align*}
	\frac{\partial}{\partial x} f(x,u) &= \frac{\partial}{\partial x} \left( Ax + B\psi(u-Cx)\right) = A - \psi'(u-Cx) BC
\end{align*}
and $\psi'(u-Cx) \in [a,b]$ for all $x$ and $u$, the proposition is proved.
\end{proof}

\end{document}